A gas X is collected over water at 17C & 750 mm pressure.If the volume of the gas colected is 50 cc,calculate the volume of a dry gas at STP .
1
Expert's answer
2018-02-15T03:19:16-0500
Vapor pressure of water at 17 °C is 14.5 mmHg
Pgas = Ptotal - Pwater
P1 = 750 – 14.5 = 735.5 mmHg = 735.5 mmHg / 760 = 0.968 atm
V1 = 50 ml = 0.05 L
t1 = 17 °C
T1 = 290 K
P2 = 1 atm
T2 = 273 K
P1V1T1 = P2V2T2
V2 = P1V1T2 / T1P2 = (0.968 atm·0.05 L·273 K) / (290 K·1 atm) = 0.0456 L = 45.6 ml